OpenStudy (anonymous):

Four-thirds times the sum of a number and 8 is 24. What is the number?

OpenStudy (anonymous):

40 26 24 10

OpenStudy (blurbendy):

4/3(x + 8) = 24 start by distributing the 4/3

OpenStudy (anonymous):

How?

OpenStudy (blurbendy):

multiply x by 4/3 and 8 by 4/3 4/3x + 32/3 = 24 now, multiply everything by 3 4x + 32 = 72 Can you solve for x now?

OpenStudy (anonymous):

x = 10

OpenStudy (blurbendy):

yup, good job